OpenStudy (anonymous):

how did Y=\frac{K}{jw + 2} become Y=\frac{K}{\sqrt{4+w^{2}}}e^{j\ominus} ???

OpenStudy (anonymous):

oopss. this is the proper how did Y=\frac{K}{jw + 2} become Y=\frac{K}{\sqrt{4+w^{2}}}e^{-j\ominus}

OpenStudy (anonymous):

you can use http://www.hostmath.com/ to translate latex format